Session: ToO observed 1Feb2010 (between session 2009/3 and 2010/1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Westerbork (Geert Kuper at Tue Feb 2 10:43:41 2010) Success :
Wb used in phased array mode. No known problems. Telescopes used: 012346789ABC

Torun (Kaz Borkowski at Tue Feb 2 10:50:54 2010) Success :
No known problems. The log, antabfas and uvflgfs files uploaded to vlbeer. Remark to JIVE: the shipment will go today and will contain two suspect disks (earlier dismounted from the pack) squeezed into the same cardbox. Appropriate note will be added to the sticker with bar code label.

Effelsberg (Uwe Bach at Tue Feb 2 11:43:04 2010) Failure :
First two scans were lost due to a disk pack failure, and then it started to snow. Effelsberg only observed for half an hour from 18 UT to 18:30 UT. The rest is noise.

Noto (P. Cassaro at Tue Feb 2 17:20:43 2010) Minor failures :
The data_valid command are "data_valid1=" in the schedule then this command give an error along the whole experiment. This could not be a serious problem, but it is there.

Onsala (Michael Lindqvist at Wed Feb 3 08:54:26 2010) Success :
No known problems.

Medicina (Andrea Orlati at Fri Feb 5 07:46:19 2010) Success :